7. What is the pH of a 0.080 M solution of the weak base pyridine, CsHsN? (K, (CsHsN) -1.7x10-*)
Answer #1

This is a base in water so, let the base be Pyridine = "B" and Pyridine H+ = HB+ the protonated base "HB+"

there are free OH- ions so, expect a basic pH

B + H2O <-> HB+ + OH-

The equilibrium Kb:

Kb = [HB+][OH-]/[B]

initially:

[HB+] = 0

[OH-] = 0

[B] = M

the change

[HB+] = x

[OH-] = x

[B] = - x

in equilibrium

[HB+] = 0 + x

[OH-] = 0 + x

[B] = M - x

Now substitute in Kb

Kb = [HB+][OH-]/[B]

Kb = x*x/(M-x)

x^2 + Kbx - M*Kb = 0

x^2 + (1.7*10^-9)x - (0.08)(1.7*10^-9) = 0

solve for x

x = 1.166*10^-5

substitute:

[HB+] = 0 + x = 1.166*10^-5 M

[OH-] = 0 + x = 1.166*10^-5M

pH = 14 + pOH = 14 + log(1.166*10^-5) = 9.066

pH = 9.066